A trader mixes three varieties of groundnuts costing Rs. 50, Rs. 20 and Rs. 30 per kg in the ratio 2 : 4 : 3 in terms of weight, and sells the mixture at Rs. 33 per kg. What percentage of profit does he make? (Ans.: 10%)
Answer by mananth(16946)

Rs. 50, Rs. 20 and Rs. 30 per kg in the ratio 2 : 4 : 3

find the cost of 1 kg mixture
50*2+20*4+30*3=270
2+4+3=9
270/9=30
cp=30
sp=33
profit = 3
profit percent = 3/30 * 100=10%